OpenStudy (iiariana):

what is the surface area of a box if it's 8x8x4 factored up by 10?

Elsa213 (elsa213):

What do you think the answer is? :o

Elsa213 (elsa213):

@iiAriana

OpenStudy (iiariana):

256 for some reason idk why

OpenStudy (anonymous):

total surface area=2(l*w)+4(l*h)

Elsa213 (elsa213):

Are there any answer choices? :o

OpenStudy (iiariana):

No I just have to just say what it is

OpenStudy (iiariana):

That's why t's so confusing

Elsa213 (elsa213):

I will let Surjithayer help you. c: Welcome to Openstudy. :D

OpenStudy (iiariana):

Ok thanks anyway

OpenStudy (anonymous):

area of base and uppersurface=2(l*W)=2(8*8)=128 area of four walls of the box=4*(8*4)=128 total area=128+128=?

OpenStudy (iiariana):

Oh. Ok thanks so much for help me

OpenStudy (iiariana):

So I was right? I said 256

OpenStudy (anonymous):

what do you mean by factored up by 10 ?

OpenStudy (iiariana):

256 factored up by 10 I think it's like *10 or divided by 10